OpenStudy (anonymous):

given the function h described h(x)=x+12 h(0)= h(-12)= h(-5)= h(12)= h(d+5)=

OpenStudy (saifoo.khan):

h(0)= 12

OpenStudy (saifoo.khan):

h(-12)= 0

OpenStudy (saifoo.khan):

h(-5)= -5 + 12 = 7

OpenStudy (saifoo.khan):

h(12)= 24

OpenStudy (saifoo.khan):

h(d+5)= (d+5) + 12 = d + 17

OpenStudy (saifoo.khan):

Here u go :)

OpenStudy (anonymous):

your amazin

OpenStudy (saifoo.khan):

hehe. Thnks!

OpenStudy (saifoo.khan):

I cant believe u guyz are taught this on Univ level. :|

OpenStudy (saifoo.khan):

Sure, Sure.

OpenStudy (anonymous):

ok find an equation of the line containing the given pair of points (4,2) and( 12,5)y=

OpenStudy (saifoo.khan):

find the slope 1st: slope = 3/8 now find the y-intercept.

OpenStudy (saifoo.khan):

y = mx + c 2 = 3 (4) + c ---- 8 c = 2 - 1.5 c = 1/2

OpenStudy (saifoo.khan):

therefore: y = mx + c y = 3/8x + 1/2 Take the LCM 8y = 3x + 4

OpenStudy (saifoo.khan):

here u go.

OpenStudy (anonymous):

so the answer is y =o

OpenStudy (saifoo.khan):

lol, no!

OpenStudy (saifoo.khan):

y = 3/8x + 1/2

OpenStudy (saifoo.khan):

this is this Answer^

OpenStudy (saifoo.khan):

lol, Yo! Bring them on!

OpenStudy (anonymous):

ok lol a i need to c that graph again